OpenStudy (anonymous):

Solve this system of eequations using substitution. Put your answe in ordered pair form, (x,y) x+2y=8 2x+y=7

OpenStudy (anonymous):

method or answer?

OpenStudy (anonymous):

Answer

OpenStudy (anonymous):

First of all you make x the subject of the formula for the first equation then you should get x=8-2y instead of x+2y=8. Then you substitute x in the second formula as 8-2y so it will become 2(8-2y)+y=7 and then you will solve it and you will get y=3. Then substitute the answer of y in any of the equations then you will get the value of x. So the answer is x=2 and y=3.

OpenStudy (anonymous):

if you only need the answer for say an on line class and you don't need to know how to do it, then do this http://www.wolframalpha.com/input/?i=x%2B2y%3D8+%2C2x%2By%3D7

OpenStudy (anonymous):

notice i just put a comma between the two equations if you need to know how to do it, then follow Faris example